In triangle ABC, AC=7, BC=12, and angle B=35 degrees. solve the triangle.
Answer by venugopalramana(3286)

In triangle ABC, AC=7, BC=12, and angle B=35 degrees. solve the triangle
AC/SIN B = BC/SIN A
7/SIN 35 = 12/SIN A
SIN A = 12*SIN 35/7=0.982838969
A = 79.4 DEG.....OR...... 100.6 DEG.
C=180-35-79.4=65.6 DEG.OR....180-35-100.6=44.4
AC/SIN B = AB/SIN C
7/SIN 35 = AB /SIN 65.6....OR.....7/SIN 35 = AB /SIN 44.4
AB = 7*SIN 65.6/SIN 35....OR..........7*SIN 44.4/SIN 35
AB = 11.12..........OR............8.54